Complete the following schedule to show the separate amounts that should be reported on the current years income statement oPrepare the journal entries for Company P at the dates indicated assuming 8,000 shares of Company T were purchased. Assume thPrepare the journal entries for Company P at the dates indicated assuming 3,200 shares of Company T were purchased. Assume thComplete the following schedule to show the separate amounts that should be reported on the current years balance sheet of CCompany T had 32,000 outstanding shares of common stock, par value $10 per share. On January 1 of the current year, Company P purchased some of Company T’s shares as a long-term investment at $23 per share. At the end of the current year, Company T reported the following: income, $48,000, and cash dividends declared during the year, $19,500. The fair value of Company T stock at the end of the current year was $20 per share.

2-a. Prepare the journal entries for Company P at the dates indicated assuming 3,200 shares of Company T were purchased. Assume the investment will be held long term.

2-b. Prepare the journal entries for Company P at the dates indicated assuming 8,000 shares of Company T were purchased. Assume the investment will be held long term.

3-a. Complete the following schedule to show the separate amounts that should be reported on the current year's balance sheet of Company P:

Required information The following information applies to the questions displayed below.] Company T had 32,000 outstanding sh3-b. Complete the following schedule to show the separate amounts that should be reported on the current year's income statement of Company P:
Answer #1

IFRS-9 :All equity investments in scope of IFRS 9 are to be measured at fair value in the statement of financial position, with value changes recognised in profit or loss, except for those equity investments for which the entity has elected to present value changes in 'other comprehensive income'. Accordingly,

2. Journal Entries :

Transaction General Journal Debit Credit
1 a (3200 shares) Investments Dr 73600 =23*3200
To Cash 73600
Cash Dr 1950
To Dividend Income 1950
Unrealized Loss Dr 9600 =3*3200
To Investments 9600
2 b(8000 shares) Investments Dr 184000 =23*8000
To Cash 184000
Investments Dr 16875 =(48000+19500)*8000/32000
Equity in Investee Earnings 16875
Unrealized Loss Dr 24000 =3*8000
To Investments 24000

3. (a) Balance Sheet :

Company P
Current Year
Dollar Amounts
3200 Shares 8000 Shares
Non Current Assets:
Investments 64000 176875
(73600-9600) (184000+16875-24000)

(b) Income Statement :

Company P
Current Year
Dollar Amounts
3200 Shares 8000 Shares
Income Statement (Other Items):
Dividend Income 1950 0
Equity in Investee Earnings 0 16875
Unrealized Loss -9600 -24000
Total -7650 -7125